OpenStudy (sarahc):

Graph the ellipse with equation x squared divided by twenty five plus y squared divided by four = 1

OpenStudy (danjs):

(x-h)/a^2 +(y-k)/b^2 = 1

OpenStudy (danjs):

x^2 / 5^2 + y^2 / 2^2 = 1

OpenStudy (danjs):

centered at the origin (0,0) +a = 5 +b = 2 semi-major and minor axis

OpenStudy (danjs):

(5,0) (-5,0) (0,2) (0,-2)